OpenStudy (jennifers731):

A white-tailed deer can sprint at speeds up to 30 miles per hour. American bison can run at speeds up to 3,520 feet per minute. Which animal is faster and by how many miles per hour? There are 5,280 feet in one mile.

OpenStudy (misssunshinexxoxo):

The american bison

OpenStudy (jennifers731):

By how many miles?

OpenStudy (anonymous):

by 40 miles

OpenStudy (misssunshinexxoxo):

3,520 x 60

OpenStudy (misssunshinexxoxo):

Please understand you need to subtract from total of the white tailed deer

OpenStudy (misssunshinexxoxo):

\[5,280 x 30 = 158,400

OpenStudy (misssunshinexxoxo):

\[211,00 - 158,000\]

OpenStudy (lilg132):

eeds up to 30 miles per hour. American bison can run at speeds up to 3,520 feet per minute. Which animal is faster and by how many miles per hour? There are 5,280 feet in one mile. white tailed deer = 30 mph american bison = 3520 feet per minute first we need to work out what both of them are in feet per hour. The white tailed deer is 30 mph. The question tells us 5280 feet in one mile. so 30 * 5280 = 158,400 feet per hour (White tailed deer) Now the american bison we already know does 3520 feet per minute so to work out how much that is per hour. (60 minutes) 3520 * 60 = 211,200 feet per hour (American bison) Now we can clearly see the american bison is than the white tailed deer in feet per hour. To work out how much faster in miles first subtract them to find out what it is in feet, 211,200 - 158,400 = 52,800 feet difference per hour but to work out in miles just divide how much feet there is in a mile. 52,800 / 5280 = 10 miles an hour faster.